OpenStudy (anonymous):

find the equation of a line passing through (-7,5) and a slope of -2

OpenStudy (anonymous):

y-y1=m(x-x1) use slope point form

OpenStudy (anonymous):

plug in the points and slope m=-2 y1=5 x1=-7 y-5=-2(x-(-7))

OpenStudy (anonymous):

that's the equation already.. but they might want you to put it y-intercept form? so solve y-5=-2(x+7) y-5=-2x-14 y=-2x-9